# Jailbreak Attacks and Defenses against Multimodal Generative Models: A Survey

### v1

@@ B.4 Output-level Defense
Output-level Defense leverages post-processing techniques to reinterpret or rephrase the generated content, aiming to neutralize potentially offensive elements while preserving the original intent.
Refiner-based defense. Refiner-based defense typically involves detecting harmful content within generated outputs, followed by employing a generative model to modify the content and eliminate identified toxicity.
$\bullet$ Image + Text → Text (IT2T) Models. MLLM-Protector [137] comprises two core components: a lightweight Harm Detector, which identifies harmful content in model outputs, and a Response Detoxifier, which transforms harmful responses into benign ones. Each component can be independently trained and seamlessly integrated with any MLLM at inference, ensuring robust safety without compromising the model’s original functionality.
@@ V Evaluation
Evaluation methods play a crucial role in providing a standardized basis for comparing various jailbreak attack and defense techniques. In this section, we first review existing evaluation datasets related to jailbreak, followed by an overview of metrics used to measure the effectiveness of jailbreak against multimodal generative models.
TABLE V: Comparison of publicly available representative evaluation datasets. For text/images from the data source, 1) Artificial: manually created by humans, 2) Website: scraped from open-world websites, 3) LLM: generated by prompting LLMs, 4) Exist: derived from existing datasets. 5) Adversarial: generated by attack methods, 6) Typography: generated by text typography into image forms, 7) Diffusion: generated by prompting T2I models.
Data Source
Dataset Task Text Image Volume Access
Any-to-Text Models
SafeBench [71] Attack LLM Typography 500 https://github.com/ThuCCSLab/FigStep
AdvBench [54] Attack LLM N/A 500 https://github.com/llm-attacks/llm-attacks
ReadTeam-2K [139] Attack LLM & Exist N/A 2,000 https://huggingface.co/datasets/JailbreakV-28K/JailBreakV-28k
HarmBench [140] Attack & Defense Unpublished N/A 320 https://github.com/centerforaisafety/HarmBench
HADES [86] Defense LLM Typography & Diffusion 750 https://github.com/AoiDragon/HADES
MM-SafetyBench [72] Defense LLM Typography & Diffusion 5,040 https://github.com/isXinLiu/MM-SafetyBench
JailBreakV-28K [139] Defense Adversarial Artificial & Exist & Diffusion 28,000 https://huggingface.co/datasets/JailbreakV-28K/JailBreakV-28k
VLGuard [112] Defense LLM Exist 3,000 https://huggingface.co/datasets/ys-zong/VLGuard
Any-to-Vision Models
NSFW-200 [76] Attack Artificial N/A 200 https://github.com/Yuchen413/text2imagesafety
MMA [19] Attack Exist & Adversarial Adversarial 1,000 https://huggingface.co/YijunYang280
VBCDE-100 [67] Attack Artificial N/A 100 https://github.com/researchcode001/Divide-and-Conquer-Attack
MPUP [141] Attack LLM N/A 1,200 https://huggingface.co/datasets/tongliuphysics/multimodalpragmatic
I2P [93] Attack & Defense Website Website 4,703 https://huggingface.co/datasets/AIML-TUDA/i2p
Unsafe Diffusion [142] Defense Artificial & Website & Exist N/A 1434 https://github.com/YitingQu/unsafe-diffusion
MACE [121] Defense Artificial Diffusion 200 https://github.com/Shilin-LU/MACE
@@ V-A Evaluation Dataset
In this part, we discuss datasets commonly used in jailbreak attacks and defenses for the multimodal generative models. These datasets are categorized into two types based on their construction intent: simulated-malicious and real-malicious datasets. Simulated-malicious datasets involve experiments conducted on pre-existing benchmarks that were not specifically designed for jailbreak scenarios. In contrast, real-malicious datasets are intentionally constructed as representative benchmarks for evaluating jailbreak vulnerabilities in multimodal generative models (see Table V).
V-A1 Simulated-malicious Dataset
@@
Due to the significant malicious potential of unsafe visual content, previous works have sought to mitigate the risk of generating genuinely harmful images in T2I jailbreak attacks. To achieve this, researchers [76, 69, 91] have typically selected specific categories from benign datasets to serve as proxies for unsafe classes, thereby avoiding the generation of real-world malicious imagery during model evaluation and testing. For instance, SneakyPrompt [76] introduces the Dog/Cat-100 dataset, which contains 100 prompts generated by GPT models describing scenarios involving dogs or cats. SneakyPrompt incorporates dog and cat concepts as part of an external safety filter. Similarly, UPAM [69] utilizes the Microsoft COCO dataset [143], a comprehensive resource containing image-text pairs for tasks such as captioning, and designates 10 specific classes (e.g., boat, bird, clock) as ”harmful” categories. Additionally, UnlearnDiffAtk [91] employs the WikiArt dataset [144], which includes fine-art paintings, along with Imagenette [145], to conduct attacks related to stylization and object manipulation.
V-A2 Real-malicious Dataset
@@
Malicious datasets typically contain explicitly harmful prompts, aligned with the usage policies of OpenAI [146] and Meta [147]. These prompts, which often involve content such as pornography, violence, and gore, are widely recognized as critical areas of safety concern due to their potential to cause discomfort and harm. Some commonly used datasets in Table V are as follows:
Image + Text → Text (IT2T) Models. Datasets in this category leverage language models or existing datasets to construct text prompts, with various options available for generating corresponding images. These options include creating blank images, adding noise, typography text onto images, or using Diffusion models to generate malicious images.
$\bullet$ SafeBench [71]. SafeBench identifies 10 safety-critical scenarios and queries GPT-4 with 50 unique malicious questions for each scenario. These questions are then rephrased into imperative sentences and converted into visual prompts through the use of typography.
@@
$\bullet$ AdvBench [54]. AdvBench prompts uncensored LLMs to generate new examples of harmful strings and behaviors by providing 5-shot demonstrations crafted by the authors. The harmful strings represent toxic or harmful topics, while harmful behaviors are framed as instructions that align with the same themes as the harmful strings.
@@
$\bullet$ RedTeam-2K [139]. RedTeam-2K dataset comprises 2,000 meticulously crafted harmful queries aimed at identifying alignment vulnerabilities within MLLMs. It incorporates queries from 16 safety policies and 8 diverse sources, including GPT Rewrite, Handcraft, GPT Generate, and several existing datasets.
@@
$\bullet$ HarmBench [140]. HarmBench encompasses 510 unique harmful behaviors, split into 400 textual behaviors and 110 multimodal behaviors, with categories specifically designed to violate laws or societal norms.
@@
Another line of jailbreak datasets often includes both standard malicious prompts and successful jailbreak prompts, specifically designed to evaluate the vulnerabilities of MLLMs. These datasets provide a rigorous basis for defense evaluation, allowing newly proposed methods to be systematically tested and validated.
$\bullet$ HADES [86]. HADES identified five safety-related scenarios and queried GPT-4 with 50 unique malicious questions for each, pairing these queries with real-world images. Additionally, the datasets include typographic images of harmful keywords and optimized images generated using stable diffusion.
$\bullet$ MM-SafetyBench [72]. MM-SafetyBench queries GPT-4 to generate multiple malicious questions for each scenario and utilizes diffusion models to produce corresponding harmful images based on key phrases extracted from the textual questions. Additionally, the generated images are combined with typography, creating a unified visual representation of the malicious content.
@@
$\bullet$ JailBreakV-28K [139]. Building on the RedTeam-2K dataset, JailBreakV-28K employs a variety of diverse jailbreak attacks to construct jailbreak prompts. These prompts are paired with four types of images: blank images, random noise images, natural images, and synthesized images generated via stable diffusion. The study covers 16 distinct themes and includes a total of 28,000 test cases, consisting of 20,000 text-based jailbreak prompts targeting LLMs and 8,000 image-based jailbreak inputs from recent MLLM jailbreak attacks.
$\bullet$ VLGuard [112]. VLGuard is dedicated to curating an instruction-following dataset that includes various harmful categories. This work gathers harmful images from existing datasets, such as Hateful Memes, and generates a total of 3,000 harmful prompts paired with these images using GPT-4V.
@@
Text → Image (T2I) Models. The datasets for T2I models are primarily focused on evaluating the generation of harmful or inappropriate content through malicious prompts.
@@
$\bullet$ NSFW-200 [76]. NSFW-200 generates 200 target prompts with malicious content using ChatGPT models.
@@
$\bullet$ MMA [19]. MMA consists of 1,000 captions with high harmfulness scores from the LAION-COCO dataset [148]. Additionally, the dataset also incorporates adversarial prompts generated using the corresponding jailbreak methods.
@@
$\bullet$ VBCDE-100 [67]. VBCDE-100 consists of 100 test cases covering five categories of violence, gore, illegal activities, discrimination, and pornographic content, with each category consisting of approximately 20 hand-curated sensitive cues.
@@
$\bullet$ MPUP [141].The MPUP dataset encompasses three high-risk scenarios: hate speech, physical harm, and fraud, with a total of 1,200 unsafe prompts. These prompts are generated by guiding GPT-4 with scenario-specific instructions, followed by manual selection and refinement.
Note that the I2P [93] dataset is widely used as a benchmark for evaluating jailbreak attacks [79, 78, 91, 75, 90] and defenses [132, 127, 128, 124, 110] in T2I models. Attack methods leverage this dataset to construct jailbreak prompts, while defense methods use these prompts to assess the robustness of proposed defense mechanisms against the generated jailbreak inputs.
@@
$\bullet$ I2P [93]. I2P encompasses a broad spectrum of inappropriate content beyond just nudity. This work gathers a total of 4,703 harmful prompts from real-world websites (i.e., Lexica), spanning seven distinct topics, and pairs them with harmful images sourced from those same websites.
In addition, other works have developed their own datasets for research, enabling a more customized approach to studying jailbreak defenses.
$\bullet$ MACE [121]. MACE offers a dataset of 200 celebrities, with their portraits generated through stable diffusion, which the Celebrity Detector identifies with notable accuracy.
@@
$\bullet$ Unsafe Diffusion [142]. Unsafe Diffusion focuses on demystifying the generation of unsafe images and hate memes from T2I models. The study collects harmful prompts from sources such as the Lexica website, and a manually created template-based dataset, along with harmless prompts from the MS COCO dataset, resulting in a collection of 1,434 prompts in total.
Text → Video (T2V) Models. As video generation models continue to evolve, the development of dedicated evaluation datasets for these models has become increasingly crucial.
@@
$\bullet$ T2VSafetyBench [20]. T2VSafetyBench is designed to evaluate the safety of T2V models, focusing on 12 critical safety aspects including pornography, violence, and discrimination. It constructs a dataset of 4,400 malicious prompts from real-world sources, GPT-generated prompts, and various jailbreak attack methods, providing a comprehensive evaluation of model safety across diverse scenarios.
@@ V-B Evaluation Method
In contrast to traditional visual-question answering datasets, generative models produce answers in an open-ended format, complicating subjective evaluation. This open-ended nature presents challenges in balancing evaluation costs with maintaining accuracy. To tackle these challenges, we have outlined diverse evaluation approaches for determining relevant metrics, as illustrated in Fig. 9. These methods are primarily classified into two categories: manual evaluation and automated evaluation.
V-B1 Manual Evaluation
@@
V-B2 Automatic Evaluation
Given the high cost and time-consuming nature of human evaluation, recent research has increasingly focused on developing automated evaluation methods. These automated approaches assess the safety of multimodal generative models by employing a range of techniques, including detector-based, GPT-based, and rule-based methods.
@@ V-B2 Automatic Evaluation
Detector-based. Detector-based approaches utilize pre-trained classifiers to automatically detect and identify harmful content within generated outputs. These classifiers are trained on large, annotated datasets that cover a range of unsafe categories, such as toxicity, violence, or explicit material.
$\bullet$ Text Detector. Toxicity detectors for detecting harmful text encompass several models, including LLama-Guard [150], LLama-Guard2 [151], Detoxify [152], GPTFuzzer [153], and Perspective API [154]. 1) LLama-Guard series models. LLama-Guard, fine-tuned on the LLaMA2 model, functions as a language model capable of performing multi-class classification while also generating binary decision scores for each harmful response. 2) Detoxify. Detoxify is working to mitigate harmful online content by identifying toxic comments. 3) GPTFuzzer. GPTFuzzer identifies jailbreak responses by fine-tuning the RoBERTa model [159] on manually curated datasets with human judgments, enabling it to assess whether a response violates safety protocols. 4) Perspective API. Perspective API employs machine learning models to identify abusive comments by scoring phrases based on their potential impact within a conversation.
@@
LLM-based. Powerful large language models like GPT-4 have increasingly been employed as safety evaluators in recent research. With carefully crafted prompts, LLMs demonstrate remarkable effectiveness in assessing whether jailbreak behaviors occur, providing nuanced judgments across various scenarios. At evaluation time, LLMs have been employed to evaluate whether the generated content from multimodal large models is semantically consistent with ground-truth annotations [72, 149, 67, 66, 20, 127, 97].
Rule-based. Rule-based approaches generally depend on predefined rules or patterns to assess the alignment between generated content and expected outputs. For instance, keywords can be used to detect the presence of essential concepts, while embedding-based matching ensures that the semantic meaning of the output aligns with ground-truth annotations. These methods offer a structured framework for evaluating both the content accuracy and semantic consistency of generated outputs.
@@
$\bullet$ Embedding-based. Other works in T2I tasks [69, 67, 65] utilize embedding-based approaches by comparing the similarity between embeddings generated by CLIP’s encoder. This method ensures that the generated images are semantically aligned with the corresponding semantic concepts by measuring the proximity of their embeddings in the latent space.

### v2

@@ B.4 Output-level Defense
Output-level Defense leverages post-processing techniques to adjust decoding strategies or rephrase the generated content, aiming to neutralize potentially offensive elements.
Decoding-based defense. Calibrating the logit distribution during decoding can steer the model’s output away from unsafe directions.
$\bullet$ Image + Text → Text (IT2T) Models. CoCA [143] uses a contrastive decoding strategy by computing the logit difference between responses with and without the safety principle, termed the safety delta. This delta is subsequently used to adjust token generation probabilities. On the other hand, IMMUNE [144] employs a safety reward function that assigns higher scores to safe responses and lower scores to unsafe ones.
Refiner-based defense. Employing a refiner (e.g., a generative model) to rephrase or edit the model’s response can also eliminate toxicity.
$\bullet$ Image + Text → Text (IT2T) Models. MLLM-Protector [145] comprises two core components: a lightweight harm detector, which identifies harmful content in model outputs, and a response detoxifier, which transforms harmful responses into benign ones. Instead of using external detoxifiers, ECSO [146] leverages the self-contained MLLM to transform malicious input images into plain texts in a query-aware manner. Safe response generation, free from image inputs, is then performed to restore the model’s intrinsic safety mechanism.
@@ V Evaluation
Evaluation methods are essential for providing a standardized basis for comparing various jailbreak attack and defense techniques. In this section, we first review existing evaluation datasets relevant to jailbreak scenarios, followed by an overview of open-ended evaluation methods and metrics.
TABLE V: Comparison of publicly available representative evaluation datasets. Collected: raw data created by humans or collected from real-world websites. Reconstructed: Data reorganized from other existing datasets. Synthesized: AI-generated data using LLM or diffusion models. Adversarial: Adversarial data generated by jailbreak attack methods. [link] directs to dataset websites.
Text Source Image Source
Dataset Venue Model Collected Reconstructed Synthesized Adversarial Collected Reconstructed Synthesized Adversarial Volume Theme
AdvBench [48] [link] [arXiv’23] I + T → T - - - - - - - 500 -
ReadTeam-2K [152] [link] [arXiv’24] I + T → T - - - - - 2,000 16
HarmBench [153] [link] [ICML’24] I + T → T - - - - - - - 510 4
Figstep [71] [link] [arXiv’23] I + T → T - - - - - - ✔ [71] 500 10
HADES [92] [link] [ECCV’24] I + T → T - - - - ✔ [92] 750 5
JailBreakV-28K [152] [link] [arXiv’24] I + T → T - - - ✔ [47, 48, 49, 50] - - 28,000 16
MM-SafetyBench [55] [link] [ECCV’24] I + T → T - - - - - ✔ [55] 5,040 13
NSFW-200 [77] [link] [SSP’24] T → I - - - - - - - 200 -
MMA [19] [link] [CVPR’24] T → I - - ✔ [19] - - - ✔ [19] 1,000 -
VBCDE [69] [link] [arXiv’23] T → I - - ✔ [69] - - - - 100 5
MPUP [154] [link] [arXiv’24] T → I - - - - - - - 1,200 4
I2P [139] [link] [CVPR’23] T → I - - - - - - 4,703 7
Unsafe Diffusion [155] [link] [CCS’23] T → I - - - - - - 1,434 -
MACE-Celebrity [130] [link] [CVPR’24] T → I - - - - - - - 1,000 -
MACE-Art [130] [link] [CVPR’24] T → I - - - - - - - 1,000 -
T2VSafetyBench [20] [link] [NeurIPS’24] T → V - ✔ [52, 53, 54] - - - - 4,400 12
@@ V-A Evaluation Dataset
In this part, we review datasets commonly used in multimodal jailbreak attacks and defenses. These datasets can be classified into two categories based on their construction intent: simulated-malicious and real-malicious datasets. Simulated-malicious datasets are typically derived from existing benign datasets that are not explicitly designed for jailbreak scenarios. In contrast, real-malicious datasets are deliberately designed to encompass a wide range of malicious topics, aiming to simulate real-world scenarios (see Table V).
A.1 Simulated-malicious Dataset
@@
Due to the significant malicious potential of unsafe visual content, previous works have sought to mitigate the risk of generating genuinely harmful images in T2I jailbreak attacks. To achieve this, researchers [77, 70, 97] have typically selected specific categories from benign datasets to serve as proxies for unsafe classes, thereby avoiding the generation of real-world malicious imagery during model evaluation and testing. For instance, SneakyPrompt [77] introduces the Dog/Cat-100 dataset, which contains 100 prompts generated by GPT models describing scenarios involving dogs or cats. Similarly, UPAM [70] utilizes the Microsoft COCO dataset [156], a comprehensive resource containing image-text pairs, and designates 10 specific classes (e.g., boat, bird, clock) as ”harmful” categories. Additionally, UnlearnDiffAtk [97] selects samples from the WikiArt [157] and Imagenette [158] datasets, to conduct attacks related to stylization and object manipulation.
A.2 Real-malicious Dataset
@@
Real-malicious datasets include prompts intended to elicit discomfort or cause harm, covering common scenarios such as pornography, violence, and gore, which are prohibited by the usage policies of OpenAI [159] and Meta [160]. Some commonly used datasets in Table V are outlined as follows:
Image + Text → Text (IT2T) Models. Existing attack methods typically utilize text-based jailbreak datasets [48, 152, 153] combined with real-world images [89, 90], AI-generated images [81, 73, 72] or adversarial images [93, 95, 87, 60] for executing multimodal jailbreaks.
@@
$\bullet$ AdvBench [48] prompts uncensored LLMs to generate two settings of harmful strings and harmful behaviors. Harmful strings encompass diverse detrimental content while harmful behaviors are framed as instructions aligned with the themes of the harmful strings.
@@
$\bullet$ RedTeam-2K [152] comprises 2,000 meticulously crafted harmful queries from 16 safety policies and 8 diverse sources, including GPT Rewrite, Handcraft, GPT Generate, and several existing datasets.
@@
$\bullet$ HarmBench [153] encompasses 510 unique harmful behaviors, split into 400 textual behaviors and 110 multimodal behaviors, which are designed to violate laws or societal norms.
@@
Another line of datasets comprises adversarial image-text pairs that can successfully jailbreak MLLMs. These datasets provide a rigorous basis for defense evaluation, enabling newly proposed defense methods to be systematically validated.
$\bullet$ Figstep [71] identifies 10 safety-critical scenarios and leverages GPT-4 to generate 50 distinct malicious questions for each scenario. These questions are subsequently transformed into imperative sentences and converted as visual prompts using typographic techniques.
$\bullet$ HADES [92] employs GPT-4 to generate 50 keywords for each harmful category and synthesize three distinct harmful instructions based on each keyword, each paired with images collected from real-world websites. Additionally, this dataset includes jailbreak images generated through the combination of typographic images, synthesized images, and adversarial noise.
@@
$\bullet$ JailBreakV-28K [152] is developed based on the RedTeam-2K [152] dataset and incorporates a diverse array of jailbreak attacks to construct jailbreak prompts. These prompts are paired with four categories of images: blank images, random noise images, natural images, and synthesized images generated using stable diffusion. The benchmark covers 16 distinct themes and encompasses a comprehensive total of 28,000 test cases.
$\bullet$ MM-SafetyBench [55] instructs GPT-4 to generate multiple malicious questions for each scenario and extract unsafe key phrases from these questions. Using the key phrases, the dataset incorporates synthesized images and typography images, creating a unified visual representation of the malicious content.
@@
Text → Image/Video (T2I/T2V) Models. These datasets primarily consist of malicious prompts designed to evaluate the generation of inappropriate content in T2I and T2V models.
@@
$\bullet$ NSFW-200 [77] includes 200 prompts with malicious content generated by using ChatGPT models.
@@
$\bullet$ MMA [19] selects 1,000 captions with high harmfulness scores from the LAION-COCO dataset [161]. Moreover, the dataset includes adversarial prompts and images generated using the corresponding jailbreak methods.
@@
$\bullet$ VBCDE [69] comprises 100 test cases covering five harmful categories of violence, gore, illegal activities, discrimination, and pornographic content, with approximately 20 sensitive prompts in each category.
@@
$\bullet$ MPUP [154] encompasses three high-risk scenarios: hate speech, physical harm, and fraud, with a total of 1,200 unsafe prompts. These prompts are generated by guiding GPT-4 with scenario-specific instructions, followed by manual selection and refinement.
@@
$\bullet$ I2P [139] comprises 4,703 harmful prompts sourced from real-world websites (i.e., Lexica), spanning seven distinct themes, and pairing them with harmful images obtained from those same websites.
Although the I2P [139] dataset is widely used for evaluating jailbreak attacks [53, 54, 97, 76, 96], a group of defense works [138, 133, 131, 128, 61] construct jailbreak prompts based on I2P to assess the robustness of proposed defense mechanisms. In addition to the I2P dataset, other datasets have been developed to enable a more customized approach to studying jailbreak defenses.
@@
$\bullet$ Unsafe Diffusion [155] focuses on demystifying the generation of unsafe images and hate memes from T2I models. The study collects harmful prompts from sources such as the Lexica website, and a manually created template-based dataset, along with harmless prompts from the MS COCO [162] dataset, resulting in a collection of 1,434 prompts in total.
$\bullet$ MACE-Celebrity [130]. To evaluate the celebrity erasure task, MACE-Celebrity consists of a database containing 200 celebrity names, each embedded within five predefined text prompts.
$\bullet$ MACE-Art [130]. To evaluate the artistic style erasure task, MACE-Art selects 200 artist names from the Image Synthesis Style Studies [163] database and applies each of them to a different set of five predefined text prompts.
@@
$\bullet$ T2VSafetyBench [20] is designed to evaluate the safety of T2V models, focusing on 12 critical safety aspects including pornography, violence, and discrimination. It constructs a dataset of 4,400 malicious prompts sourced from existing datasets, GPT-generated prompts, and jailbreak attack-based prompts, providing a comprehensive evaluation of model safety across diverse scenarios.
@@ V-B Evaluation Method
In contrast to traditional visual-question answering datasets [164, 165], generative models produce answers in an open-ended format, complicating subjective evaluation. This open-ended nature presents challenges in balancing evaluation costs with maintaining accuracy. To tackle these challenges, we have outlined diverse evaluation methods for determining relevant metrics, as illustrated in Fig. 9. These methods are primarily classified into two categories: manual evaluation and automated evaluation.
B.1 Manual Evaluation
@@
B.2 Automatic Evaluation
Given the high cost and time-consuming nature of human evaluation, recent works [148, 155, 55, 166, 90] have increasingly focused on developing automated methods to evaluate the toxicity of generated content across different modalities. These approaches can be categorized into three types: detector-based evaluation, GPT-based evaluation, and rule-based evaluation.
@@ B.2 Automatic Evaluation
Detector-based Evaluation. Detector-based evaluation utilizes pre-trained classifiers to automatically detect toxic content within generated outputs. These classifiers are trained on extensive, annotated datasets that encompass a broad spectrum of unsafe categories, including toxicity, violence, or explicit material.
$\bullet$ Text Detector. Toxicity detectors for detecting harmful text encompass several models, including LLama-Guard [148], LLama-Guard2 [149], Detoxify [167], GPTFuzzer [168], and Perspective API [169]. 1) LLama-Guard series models. LLama-Guard and LLama-Guard2 are fine-tuned on the LLaMA-based [172] models, enabling them to perform multi-class classification and generate binary decision scores for harmful responses. 2) Detoxify. Detoxify is working to mitigate harmful online content by identifying toxic comments. 3) GPTFuzzer. GPTFuzzer identifies harmful responses by fine-tuning the RoBERTa model [173] on manually curated datasets with human judgments. 4) Perspective API. Perspective API employs machine learning models to identify abusive comments by scoring phrases based on their potential impact within a conversation.
@@
LLM-based Evaluation. Powerful LLMs especially for GPT-4, have increasingly been employed as safety evaluators in recent research [55, 166, 69, 82, 20, 133, 104]. Unlike specialized detectors constrained by specific themes, LLMs exhibit exceptional effectiveness in identifying jailbreak behaviors across diverse scenarios. Specifically, LLMs are employed to perform nuanced semantic analysis to assess the consistency between generated content and ground-truth annotations.
Rule-based Evaluation. Rule-based evaluation typically depends on predefined rules or patterns to assess the alignment between generated content and expected outputs. For instance, keywords can be used to identify the presence of essential concepts, while embedding-based matching ensures that the semantic meaning of the output aligns with ground-truth annotations.
@@
$\bullet$ Embedding-based. Embedding-based approaches in T2I tasks [70, 69, 67] aim to evaluate the semantic quality of generated images. By extracting high-dimensional vectors from encoders such as the CLIP model, these methods can compute their cosine similarity score to assess semantic alignment between the generated images and target harmfulness.